OpenStudy (anonymous):

which of the following are the solutions to the equation below? Check all that apply. (2x-3)^2=1 a.-1 b.3 c.-2 d.1/3 e.2 f.1

OpenStudy (anonymous):

Substitute all the values into x and see if you get 1, if you did, then that value of x is a solution to the equation

OpenStudy (anonymous):

B. and D. are out of the list

OpenStudy (anonymous):

Still got more out of list

OpenStudy (anonymous):

No

OpenStudy (mathmate):

Since there are multiple solutions, you are better off finding the solutions, instead of testing all 5 choices. \( (2x-3)^2 =1 \\ => \\ (2x-3)=\pm 1 \) So solve for \( 2x-3=1\\ and\\ (2x-3)=-2 \) to find your answer.

OpenStudy (mathmate):

Sorry, last one should read: \( (2x-3)=-1\)

OpenStudy (anonymous):

No

OpenStudy (anonymous):

c nd e??

OpenStudy (mathmate):

Did you solve for x in the two equations I posted earlier? What did you get?

OpenStudy (anonymous):

2,-2

OpenStudy (mathmate):

Can you check your second answer?

OpenStudy (anonymous):

wait it -1 got confussed

OpenStudy (anonymous):

its a nd e

OpenStudy (mathmate):

Sure it's -1?

OpenStudy (mathmate):

Are you sure the second answer is -1?

OpenStudy (mathmate):

Solve \((2x−3)=−1\)

OpenStudy (anonymous):

for the second equation there both - so its +1

OpenStudy (mathmate):

So what are the solutions of the equations?

OpenStudy (anonymous):

i mean -2

OpenStudy (anonymous):

ohhhhh d nd e

OpenStudy (anonymous):

No

OpenStudy (mathmate):

For the second equation: \((2x−3)=−1\\ 2x=-1+3=2\\ x=1\) Do you agree?

OpenStudy (anonymous):

isnt it -1

OpenStudy (mathmate):

Show me how you got -1, please.

OpenStudy (anonymous):

-+=- right

OpenStudy (mathmate):

@romanortiz65 you still there?

OpenStudy (anonymous):

yeah

OpenStudy (mathmate):

So can you show me how did you get x=-1 from \((2x−3)=−1\) This will help you avoid future errors.

OpenStudy (anonymous):

2x-3=-1 +3 +3 2x=-2 x=-1

OpenStudy (anonymous):

-1 + 3 = 3 - 1 = 2

OpenStudy (mathmate):

On the right hand side, -1+3 equals -2. I will show this on the number line: 

OpenStudy (mathmate):

Sorry again, -1+3=+2

OpenStudy (mathmate):

So what are the two solutions that we have found?

OpenStudy (anonymous):

-1, 2

OpenStudy (anonymous):

A,E

OpenStudy (anonymous):

No

OpenStudy (anonymous):

what the heck im confffused

OpenStudy (anonymous):

e nd f

OpenStudy (anonymous):

its okay... slow down

OpenStudy (anonymous):

There you go!

OpenStudy (anonymous):

\[2x = 4\]and \[2x - 3= -1\]Thats how you get E and F